I have 1.01 of a solution that is 0.100M Fe(NO3)2 and 0.200M Cu(NO3)2​. I want to seperate the metal ions. I gradually add a solution that is 0.100M in K3PO4, 0.100 M in KOH, and 0.100 M in K2CO3.What volume of the solution do I need to add to get the first metal to begin to precipitate?
Answer #1

We need to ascertain what metal salts can possibly precipitate out from the solution.

We have the following options.

Fe3(PO4)2 (s) <=====> 3 Fe2+ (aq) + 2 PO43- (aq); Ksp = 1.0*10-36

Fe(OH)2 (s) <=====> Fe2+ (aq) + 2 OH- (aq); Ksp = 7.9*10-16

FeCO3 (s) <======> Fe2+ (aq) + CO32- (aq) ; Ksp = 2.1*10-11

Cu3(PO4)2 (s) <=====> 3 Cu2+ (aq) + 2 PO43- (aq); Ksp = 1.40*10-37

Cu(OH)2 (s) <=====> Cu2+ (aq) + 2 OH- (aq); Ksp = 4.8*10-20

CuCO3 (s) <=====> Cu2+ (aq) + CO32- (aq); Ksp = 1.40*10-10

We see that both the metal ions are divalent and form salts that solubilize in the same way. Hence, we can compare the Ksp values to predict which metal ion will precipitate out first. Since Fe3(PO4)2 and Cu3(PO4)2 have comparable Ksp values, either one of them will precipitate out first.

Find out the concentration of PO43- required to precipitate out both.

Fe3(PO4)2: Ksp = [Fe2+]3[PO43-]2

===> 1.0*10-36 = (0.100)3*[PO43-]2

===> [PO43-]2 = 1.0*10-33

===> [PO43-] = 3.162*10-17

The concentration of PO43- required to precipitate out Fe3(PO4)2 is 3.162*10-17 M. Since the concentration of PO43- is extremely low, we can assume the total volume of the solution to stay constant at 1.0 L.

Use the dilution equation to find out the volume of K3PO4 required.

M1*V1 = M2*V2

===> (1.0 L)*(3.162*10-17 M) = (0.100 M)*V2

===> V2 = 3.162*10-16 L (ans).

Cu3(PO4)2: Ksp = [Cu2+]3[PO43-]2

===> 1.40*10-37 = (0.200)3*[PO43-]2

===> [PO43-]2 = 1.75*10-35

===> [PO43-] = 4.183*10-18

The concentration of PO43- required to precipitate out Cu3(PO4)2 is 4.183*10-18 M. Since the concentration of PO43- is extremely low, we can assume the total volume of the solution to stay constant at 1.0 L.

Use the dilution equation to find out the volume of K3PO4 required.

M1*V1 = M2*V2

===> (1.0 L)*(4.183*10-18 M) = (0.100 M)*V2

===> V2 = 4.183*10-17 L (ans).

Since it requires a lower volume of PO43- to precipitate out Cu3(PO4)2, hence it will take 4.183*10-17 L of the solution to precipitate out the first salt.